# references.bib

### v1

@@ 3 Experimental Setup
To recapitulate, we structure our study into two experiments. In the first experiment (subsection 4.1), we aim to see if a language model can generally predict the gender of a character given different amounts of speech and different attribution methods. This experiments adopts a two-factor design. The first factor we vary is the input length, and the second one is the aggregation strategy. The three input levels are 1) Character level, in which we consider all the lines that a character speaks throughout the course of a play, 2) Scene level, in which we consider all of the lines that a character speaks in a given scene, and 3) Utterance level, in which we consider only unique lines (or utterances) spoken by a character. The aggregation strategies are: 1) None, 2) majority label, and 3) geometric mean, as discussed in Section 2.
The second experiment (subsection 4.2) focuses specifically on cross-dressing characters which exhibit a broad range of social roles in Calderóns plays. In his own words, spoken by Rosaura in La vida es sueño:
Tres veces son las que ya me admiras, tres las que ignoras quién soy, pues las tres me has visto en diverso traje y forma. La primera me creíste varón, en la rigurosa prisión, donde fue tu vida de mis desdichas lisonja. La segunda me admiraste mujer, cuando fue la pompa de tu majestad un sueño, una fantasma, una sombra. La tercera es hoy, que siendo monstruo de una especie y otra, entre galas de mujer armas de varón me adornan.
Three times you have admired me, three times you do not know who I am, for three times you have seen me in different costume and form. The first time you believed me to be a man, in the rigorous prison, where your life was a flattery of my misfortunes. The second you admired me woman, when the pomp of your majesty was a dream, a phantom, a shadow. The third is today, that being a monster of one kind and another, among the finery of a woman, the weapons of a man adorn me.
Investigating specifically these characters tells us to what extent the model potentially only picks up ’prototypical’ scenes of characters, and to what extent it is able to pick up more fine-grained change of roles across scenes within characters.
@@
We base our study on 109 of Calderón’s plays that have previously been classified as non-religious plays of various genres, jointly known as ’comedias’, that are available on the CalDraCor part of the DraCor project [Fischer2020ProgrammableCI]. [^3] https://dracor.org/cal. These are orthographically modernized and normalized digital versions of the plays in TEI-XML format. DraCor segments texts by utterances in the style of classical play scripts, with lines labeled with the speaker. The plays are already annotated by act number but lack formal scene division. We enriched the plays with scene divisions, determined by stage directions that indicate a character leaving or entering the stage, to obtain a more fine-grained segmentation.
In the cast list of each play, all characters are annotated with role information, including ’male’ and ’female’ labels. These annotations are based on examination of the texts done by literary scholars as part of the DraCor project. Note that the intention of this paper is not to perpetuate the idea of a gender binary, nor to make assertions about gender identity or presentation.
In order to define a simple binary classification task, we leave out a small number of characters (such as ’musicos’ or ’coro’ (musicians, choir)) which do not have a defined gender in the corpus. After also removing minor characters who speak less than 30 words — presumably not enough for reliable classification — we are left with a total of 1515 characters (1021 = Male, 494 = Female): We have roughly a 2:1 male–female character ratio. Males speak a mean of 1219 tokens (min = 31, max = 7166) and females on average 1130 tokens (min = 32, max = 5388). On this basis, we assume that text length is not a major confounding variable for the male/female distinction.
We create an 80%/10%/10% train/test/validation split at the character level, ensuring that each character occurs only n in one of the partitions (1207/152/152 characters). Since we only provide the character speech to the classifier, this means that the test and validation partitions are essentially unseen.
@@
To avoid the lexical overfitting problem discussed above, we masked all names of characters that appear in the plays, there were 751 unique names masked with the (pseudo-)token [NAME]. We also masked all tokens that appeared in only one play with a [MASK] token. We also considered removing signals of grammatical gender such as self-predications of the speaker with gender-specific nouns or gendered adjective endings. However, automating such a filter would require some amount of semantic analysis to recognize which predicative constructions apply to the speaker themselves, arguably a non-trivial process that runs the risk of introducing inconsistencies. Also, these cases appear sufficiently rarely in the corpus to not have a major effect, as we demonstrate below.
@@
Following the argument from the previous section, we build our classifiers by fine-tuning a pre-trained Language Model on our drama data. We tested the model performance on a series of multilingual BERT, DistilBERT, and RoBERTa models. We found the best performance for the classification task was achieved when using BETO [CaneteCFP2020], a BERT base model pre-trained on Spanish data, namely 3B token from Spanish Wiki, Europarl, and other corpora. This LM provides a suitable basis since the comedias are modernized and normalized, a discussed above.
Note that BETO, like all BERT-style models, carries out internal tokenization using a variant of byte pair encoding [sennrich-etal-2016-neural]. This method has the goal of limiting the vocabulary of tokens in text. Its consequence is that frequent words tend to remain separate tokens, while infrequent words are split into subword units, minimally individual characters. By convention, subword tokens that start inside original words are prefixed with ## (cf. Figure 2). Also, BETO has a maximum context window size of 512 tokens, meaning that for longer texts, only the first 512 tokens are taken into account in any processing. This amount of text is generally sufficient for the model to make predictions. Details on model training can be found in subsection 6.1.
@@
We evaluate both the individual predictions and the aggregated predictions for characters quantitatively with precision, recall, and F1 score, macro-averaged between the ’male’ and ’female’ classes. We use macro averaging because, while the male class is over-represented compared to the female class, we are equally interested in male and female predictions.
For the qualitative evaluation, we use an explainability library that provides an implementation of integrated gradients (Transformers Interpret Sequence Classification Explainer) [^4] https://github.com/cdpierse/transformers-interpret. It gives us the contribution of each token towards the overall prediction of each test input sequence.
@@ 4.1 Experiment 1: General Gender Classification
Table 1 shows the main results for gender classification at the three input levels without any aggregation. The model performs best at the character level (F1 = 0.83), followed by the scene (F1 = 0.70) and utterance (F1 = 0.68) levels. This is not very surprising, given that the character level provides the model with the longest, and thus most informative, input texts. It it interesting, though, that the performance of the model drops by 15 points F1 score between the character and the utterance level, while the performance of the baseline only drops by 3 points F1 score. The outcome of this analysis, therefore, seems to be: more text is better.
@@
Character 0.33 0.50 0.38 0.83 0.82 0.83
Scene 0.33 0.50 0.40 0.75 0.69 0.70
Utterance 0.34 0.50 0.41 0.71 0.67 0.68
@@
To further explore this pattern, Table 2 shows F1 scores for our gender classifiers on the four quartiles of characters in terms of numbers of words spoken. The results again demonstrate a clear improvement for longer texts of 15-20 points F1 score between the first and the forth quartile. We take this patterns as evidence that, as conjectured above, many short utterances are not informative enough to permit reliable gender determination of the speaker.
@@
Utterances 0.63 0.65 0.69 0.78
Table 2: F1 for test set characters divided into quartiles by number of words spoken
@@
Next, Table 3 presents the performance of the models when classification is first carried out at the utterance or scene level and then aggregated to the character level. For this reason the numbers in Table 3 are comparable to the ’Character’ row of Table 1.
Comparing the two aggregation methods, we observe that majority vote performs rather badly, leading to a substantial loss in performance compared to the predictions without aggregation in Table 1. In contrast, combination by geometric mean performs well, and predictions at the character level almost reach the performance of the model that consider the complete text of a character.
This pattern meshes well with our interpretation above: when we use the probability estimates of the model, uninformative utterances have less impact on the overall prediction, and performance improves. All in all, we believe that the aggregated scene-level model is the most interesting among the ones we consider, since it is almost as good as the character-level model when predictions at the character level are evaluated, but also provides a separate prediction for each character in each scene, making it possible to pick up changes in gender roles over the course of a play – exactly the phenomenon that we are going to investigate below in Experiment 2. Thus, our second analysis yields a somewhat more nuanced picture: more text is better, but the effect can be mitigated by informed aggregation.
@@
Scene 0.51 0.51 0.49 0.81 0.81 0.81
Utterance 0.33 0.31 0.32 0.75 0.79 0.76
@@
Character 0.87 0.70 - - - -
Scene 0.83 0.58 0.72 0.25 0.87 0.744
Utterance 0.82 0.54 0.51 0.12 0.82 0.71
Table 4: F1 Scores by Gender
@@
Table 4 shows the results of the non-baseline models broken down by gender; best results for each metric and each gender are boldfaced. The results for male characters are generally better than for female characters, which is expected, given the higher number of male characters. That said, there is a striking difference in the effect of the aggregation strategy between genders: while male results show only a rather small change with 5 points between the Character and Utterance levels, female results drastically improved at the higher input levels (+.04 at the scene level and +16 at the Character level. When taking into account just the majority label for each character, the performance drops, especially for the female characters, at both the scene and utterance level. When calculating the geometric mean, the performance at the scene and utterance level for both male and female characters increases, but especially for female characters. Again, this goes well with our general observation that some sentences are highly gendered while others are uninformative as far as the model is concerned. Thus, the geometric mean-based aggregation can be seen as a strategy to reduce gender bias, according to its definition as performance difference between genders [rudinger-etal-2018-gender].
@@
To further validate this interpretation, we examined both the 10 correct and 10 incorrect characters with the highest confidence according to the model. It was assumed that the gender annotations found in the DraCor corpus were gold standard. However, through this examination we identified several annotation mistakes in the corpus. Among the 10 most confident ’incorrect’ male predictions, were two female characters incorrectly annotated as male (Palas and Glauca). Likewise, among the 10 most confident ’incorrect’ female annotations, we found an erroneously annotated character (Simón).
Attribution Model
@@
From the quantitative model evaluation, it is clear that the model is able to predict gender of characters in Calderón’s plays and even catch some annotation mistakes. However, it is not clear yet to what extent the model drew on features of the input that actually have to do with gender portrayal, despite our efforts to remove confounder features (cf. the discussion on lexical overfitting in section 2). By using an attribution model, we now analyze concretely on the basis of which tokens the model makes its classification decisions. In the following analysis, we list BETO tokens together with their integrated gradient scores, where negative scores indicate a tendency towards female speech, and positive scores a tendency towards male speech. Values closer to 0 are gender neutral, according to the model.
In order to determine if there were clear themes in the tokens that the model judged as being more masculine or more feminine, we first examined the averaged attribution scores of each token, calculated by averaging the score of the token across the validation set. Table 5 lists the 20 tokens deemed to be the most feminine and masculine, respectively. [^5] For convenience, the Appendix provides a translation of this table into English as Table 8. It is important to note that in each context, the attribution score is different for each context in which the token appears.
@@
(’nosotras’, 0.7763) 2 (’ton’, -0.8746) 1
(’##jote’, 0.6984) 1 (’cansada’, -0.8000) 1
(’Duque’, 0.6460) 1 (’Ami’, -0.6729) 1
(’Emperador’, 0.6105) 2 (’público’, -0.6722) 1
(’Majestad’, 0.6073) 3 (’desesperada’, -0.6178) 2
(’escu’, 0.6008) 1 (’contenta’, -0.5764) 1
(’tropas’, 0.5860) 1 (’Todas’, -0.4795) 16
(’infantería’, 0.5839) 1 (’esperaré’, -0.4431) 1
(’##estre’, 0.5824) 1 (’beso’, -0.4076) 1
(’encerrado’, 0.5733) 1 (’minuto’, -0.3882) 1
(’Alá’, 0.5613) 1 (’amigas’, -0.3841) 2
(’Gobernador’, 0.5448) 2 (’damas’, -0.3609) 4
(’Nevada’, 0.5178) 1 (’##ísimas’, -0.3163) 1
(’Fur’, 0.5095) 5 (’Dis’, -0.3154) 2
(’camarada’, 0.5076) 2 (’##acas’, -0.3066) 1
(’alojamiento’, 0.4962) 3 (’discursos’, -0.2966) 1
(’Fox’, 0.4954) 1 (’##món’, -0.2954) 1
(’soldado’, 0.4922) 2 (’cuerpo’, -0.2901 ) 1
(’torneo’, 0.4919) 1 (’comedia’, -0.2895) 1
(’criados’, 0.4915) 1 (’celosa’, -0.2844) 1
@@
We can see several patterns in the attributions of tokens made by the model. Included in the most feminine tokens are several word endings or words with feminine endings, (cansada, desesperada, contenta, celosa, todas, ##ísims, ##acas). These grammatically grounded indicators of female characters indicate that in Calderón’s plays, female characters either refer to themselves or other women frequently.
@@
Moving on to a more semantic level, we see that some of the most feminine words have some semantic connection to interpersonal relationships between characters(amigas, beso, discursos, cuerpo, celosa). This indicates that, broadly throughout Calderón’s works, female characters are more likely to discuss themes of love and relationships than male characters. Other of these words (cansada, desesperada, contenta, celosa) also demonstrate a depiction of women as emotional - the women in these plays are on one hand able to discuss their emotional inner worlds, but on the other hand depicted as driven or controlled by their emotions [drouet2017crimen]. This observation ties to larger cultural concepts about women as emotional beings and men as logical. Here we can see that Calderón’s work somewhat conforms to these norms, but we will see later some evidence of the subversion of this trope in 4.2.
@@
In contrast, the male list contains several titles (Majestad, Emperador, Duque, Gobernador), which indicates the same pattern of more within-gender than across-gender interactions for male characters as well. Furthermore, these words indicates the importance of status in male-male interactions, arguably closely related to the topic of honor which was established as important across many words by Calderónian scholarship. [^6] Note however the absence of words for honor itself (honor, honria), which makes sense, since honor is relevant for both men and women, even if in different contexts. Corresponding to the female indicators for conversations about relationships, the male list shows several words semantically related to combat (tropas, infantería, torneo, soldado), apparently the prototypical male topic. As previously mentioned, honor is of primary importance in Calderón’s works. In addition to the fact that war is a typically male endeavor, in the Siglo de oro, male characters frequently use violence as a means to gain or restore honor [honig1961seizures]. It is therefore unsurprising that these themes feature prominently as ’masculine’ words.
@@
Curiously, nosotras is the token most attributed as male, however, it is not spoken by a male character in the corpus. It appears in the validation set 8 times, and of the 8 characters who do speak this token, 7 of them are correctly classified as female characters. Upon further investigation, it was found that nosotras appears in the training corpus, spoken by a female character that is mislabeled as male. Its listing as a female token is therefore the result of an annotation mistake. We will address these annotation mistakes in future work.
@@
We also analyzed a list of top-40 male and female tokens in terms of formal linguistic properties, given that word classes or word types can also be used for stylistic gender portrayal [caliskan2022gender]. We found that the most female tokens included significantly more word fragments (15), compared to the male tokens (10). Conversely, the male tokens included more nouns (24) compared to the female tokens (14). The female tokens also included slightly more adjectives (5) compared to male (3). Both used an equal number of verbs. In conclusion, there do not appear to be major differences at the level of word classes – if anything, female characters tend to use less frequent words that tend to get split up into subword tokens.
@@
Overall, we believe that this inspection of the features used by the model indicates that the model makes its calls reasonably based on words in the input. They indicate that Calderón’s gender portrayal – within the space of features considered by the model – has two main components: (a) at the interaction level, women and men both prefer interlocutors of the same gender; (b) at the content level, women are more concerned with the domestic sphere and men the military sphere. In contrast, there do not seem to be a major formal linguistic differences between the two genders.
@@
Note that we only inspected the top tokens associated with either gender, and therefore this analysis cannot make any strong claim as to completeness. It is indeed likely that other topics arise from a further analysis of such tokens. Furthermore, the masking procedure we carried out during processing can be repeated after this type of analysis: if, for example, the analyst decides that tokens indicating patterns of interaction are not interesting in terms of gender portrayal, these tokens can be masked, and the model training repeated. The results is a model which should have the ’next best’ set of features among its top-ranking gender cues.

### v2

@@ 3 Experimental Setup
To recapitulate, we structure our study into two experiments. In the first experiment (subsection 4.1), we aim to see if a language model can generally predict the gender of a character given different amounts of speech and different attribution methods. This experiments adopts a two-factor design. The first factor we vary is the input length, and the second one is the aggregation strategy.
The three input levels are 1) character level, in which we consider all the lines that a character speaks throughout the course of a play as one data-point, 2) scene level, in which we consider all of these lines, but divided by the scene barriers so that each scene is predicted separately, and 3) utterance level, in which we consider only unique lines (or utterances) spoken by a character. The aggregation strategies are: 1) none, 2) majority label, and 3) geometric mean, as discussed in section 2. This experiment targets the maximum number of characters from Calderón’s plays for whom we can obtain reliable digitized text (see below for details).
The second experiment (subsection 4.2) focuses specifically on cross-dressing characters, who exhibit a broad range of social roles in Calderóns plays. In his own words, spoken by Rosaura in La vida es sueño:
Tres veces son las que ya / me admiras, tres las que ignoras / quién soy, pues las tres me has visto / en diverso traje y forma. / La primera me creíste / varón, en la rigurosa / prisión, donde fue tu vida / de mis desdichas lisonja./ La segunda me admiraste / mujer, cuando fue la pompa / de tu majestad un sueño,/ una fantasma, una sombra./ La tercera es hoy, que siendo / monstruo de una especie y otra,/ entre galas de mujer / armas de varón me adornan.
Three times you have / admired me, three times you do not know / who I am, for three times you have seen me / in different costume and form./ The first time you believed me to be / a man, in the rigorous / prison, where your life was / a flattery of my misfortunes./ The second you admired me / woman, when the pomp / of your majesty was a dream,/ a phantom, a shadow./ The third is today, that being/ a monster of one kind and another,/ among the finery of a woman,/ the weapons of a man adorn me.
Investigating specifically these characters tells us to what extent the model potentially only picks up ‘prototypical’ scenes of characters, and to what extent it is able to pick up more fine-grained change of roles across scenes within characters.
@@
We base our study on all 109 of Calderón’s plays that have previously been classified as non-religious plays of various genres, jointly known as ‘comedias’, which are available in the CalDraCor part of the DraCor project ([fischer_2020]). [^3] See: https://dracor.org/cal. These are orthographically modernized and normalized digital versions of the plays in XML-TEI format. DraCor segments texts by utterances in the style of classical play scripts, with lines labeled with the speaker. The plays are already annotated by act number but lack formal scene division. We enrich all plays with scene divisions, determined by stage directions that indicate a character leaving or entering the stage, to obtain a more fine-grained segmentation.
In the cast list of each play, all characters are annotated with role information, including ‘male’ and ‘female’ labels. These annotations are based on examination of the texts done by literary scholars as part of the DraCor project. Note that the intention of this paper is not to perpetuate the idea of a gender binary, nor to make assertions about gender identity.
In order to define a simple binary classification task, we leave out a small number of characters who do not have a defined gender in the corpus, typically collectives (such as “musicos” or “coro” (musicians, choir). After also removing minor characters who speak less than 30 words – presumably not enough for reliable classification – we are left with a total of 1,515 characters (1,021 = Male, 494 = Female): We have roughly a 2:1 male/female character ratio. Males speak a mean of 1,219 tokens (min = 31, max = 7,166) and females on average 1,130 tokens (min = 32, max = 5,388). On this basis, we assume that text length is not a major confounding variable for the male/female distinction.
We create an 80%/10%/10% train/test/validation split at the character level, ensuring that each character occurs only in one of the partitions (1,207/152/152 characters). Since we only provide the character speech to the classifier, this means that the test and validation partitions are unseen.
@@
To avoid the lexical overfitting problem discussed above, we mask all names of characters that appear in the plays (751 unique names) with the token ‘[NAME]’. We also mask all tokens that appeared in only one play with a ‘[MASK]’ token. We also considered removing signals of grammatical gender such as self-predications of the speaker with gender-specific nouns or gendered suffixes (such as “desesperada” above). However, automating such a filter would require substantial linguistic analysis to recognize which predications apply to the speaker themselves, arguably a non-trivial process that runs the risk of introducing inconsistencies. Also, these cases appear sufficiently rarely in the corpus to not have a major effect, as we demonstrate below.
@@
Following the argument from the previous section, we build our classifiers by fine-tuning a pre-trained Language Model on our drama data. We tested the model performance on a series of multilingual BERT, DistilBERT, and RoBERTa models. We found the best performance for the classification task was achieved when using BETO ([canete_2020]), a BERT base model pre-trained on Spanish data, namely 3B token from Spanish Wiki, Europarl, and other corpora. We adopt this LM for our experiments. It provides a suitable basis given that the dramas we work with are modernized and normalized.
Note that BETO, like all BERT-style models, carries out internal tokenization using a variant of byte pair encoding ([sennrich_2016]). This method has the goal of limiting the vocabulary of tokens in text. Its consequence is that frequent words tend to remain separate tokens, while infrequent words are split into subword units, minimally individual characters. By convention, subword tokens that start inside original words are prefixed with ‘##’ (Figure 2). Also, BETO has a maximum context window size of 512 tokens, meaning that for longer texts, only the first 512 tokens are taken into account in any processing. This amount of text is generally sufficient for the model to make predictions. Details on model training can be found in subsection 6.1.
@@
We evaluate both the individual predictions and the aggregated predictions for characters quantitatively with precision, recall, and F1 score, macro-averaged between the ‘male’ and ‘female’ classes. We use macro averaging because, while the male class is over-represented compared to the female class, we are equally interested in male and female predictions.
For the qualitative evaluation, we use a library that provides an implementation of the integrated gradients approach (Transformers Interpret Sequence Classification Explainer). [^4] See: https://github.com/cdpierse/transformers-interpret. It yields the contribution of each token towards the overall class prediction of each test input sequence.
@@ 4.1 Experiment 1: General Gender Classification
Table 1 shows the main results for gender classification at the three input levels without any aggregation. The model performs best at the character level (F1 = 0.83), followed by the scene (F1 = 0.70) and utterance (F1 = 0.68) levels. This is not very surprising, given that the character level provides the model with the longest, and thus most informative, input texts. It is interesting, though, that the performance of the model drops by 15 points F1 score between the character and the utterance level while there is hardly any difference between the scene level and the utterance level. The baseline results change hardly between levels, indicating that this effect does not reflect a change in the difficulty of the task. We therefore draw as our first conclusion that reliable gender identification requires more information than is typically available from a single utterance or even scene.
@@
Character per drama 0.33 0.50 0.38 0.83 0.82 0.83
Character per scene 0.33 0.50 0.40 0.75 0.69 0.70
Character utterance 0.34 0.50 0.41 0.71 0.67 0.68
@@
F1 for character utterance 0.63 0.65 0.69 0.78
Table 2: F1 for test set characters divided into quartiles by number of words spoken.
To explore the failure modes, Table 2 shows F1 scores for our gender classifiers at the scene and utterance levels. We split these numbers into four quartiles of characters in terms of numbers of words spoken. The results again demonstrate a clear improvement for longer texts of 15-20 points F1 score between the first and the forth quartile. Evidently, many short utterances are not informative enough to permit reliable gender determination of the speaker.
@@
Next, Table 3 presents the performance of the models when classification is first carried out at the utterance or scene level and then aggregated to the character level. The numbers in Table 3 are comparable to the ‘Character per drama’ row of Table 1.
@@
Character per scene 0.51 0.51 0.49 0.81 0.81 0.81
Character utterance 0.33 0.31 0.32 0.75 0.79 0.76
@@
Comparing the two aggregation methods, we observe that the majority vote performs rather badly, leading to a substantial loss in performance compared to the predictions without aggregation in Table 1. In contrast, the combination by geometric mean performs well, and aggregated predictions at the scene level almost reach the performance of the model that considers the complete text of a character.
This pattern meshes well with our interpretation above: When we use the probability estimates of the model, uninformative utterances have less impact on the overall prediction, and performance improves. All in all, we believe that the aggregated scene-level model is the most interesting among the ones we consider, since it is almost as good as the character-level model when predictions at the character level are evaluated, but also provides a separate prediction for each character in each scene, making it possible to pick up changes in gender roles over the course of a play – exactly the phenomenon that we are going to investigate below in Experiment 2 (subsection 4.2). Thus, our second analysis yields a somewhat more nuanced picture: Gender classification requires a fair amount of text, but the effect can be mitigated by informed aggregation.
@@
Character per drama 0.87 0.70 - - - -
Character per scene 0.83 0.58 0.72 0.25 0.87 0.74
Character utterance 0.82 0.54 0.51 0.12 0.82 0.71
Table 4: Model F1-Scores comparing performance on predictions of male and female characters.
@@
Table 4 shows the results of the non-baseline models broken down by gender; best results for each metric and each gender are boldfaced. The results for male characters are generally better than for female characters, which is expected, given the higher number of male characters. That said, there is a striking difference in the effect of the aggregation strategy between genders: While results for male characters show only a rather small change of 5 points between the character and utterance levels, results for female drastically improve at the higher input levels (+.4 at the scene level and +16 at the drama level). When taking into account just the majority label for each character, the performance drops, especially for the female characters, at both the scene and utterance levels. When calculating the geometric mean, the performance at the scene and utterance level for both male and female characters increases again, in particular for female characters. Again, this goes well with our general observation that some sentences are highly gendered while others are uninformative as far as the model is concerned. Thus, the geometric mean-based aggregation can be seen as a strategy to reduce gender bias, according to its definition as performance difference between genders ([rudinger_2018]).
@@
To further validate this interpretation, we examined both the 10 correct and 10 incorrect characters with the highest confidence according to the model. We initially assumed that the gender annotations found in the DraCor corpus were gold standard. However, through this examination we identified several annotation mistakes in the corpus. Among the 10 most confident ‘incorrect’ male predictions, were two female characters incorrectly annotated as male (Palas and Glauca). Likewise, among the 10 most confident ‘incorrect’ female annotations, we found an erroneously annotated character (Simón).
Attribution Model.
@@
From the quantitative model evaluation, it is clear that the model is able to predict the gender of characters in Calderón’s plays and even identify some annotation mistakes. However, it is not clear yet to what extent the model drew on features of the input that actually have to do with gender portrayal, despite our efforts to remove confounder features (see the discussion on lexical overfitting in section 2). By using an attribution model, we now analyze concretely on the basis of which tokens the model makes its classification decisions. In the following analysis, we list BETO tokens together with their integrated gradient scores, where negative scores indicate a tendency towards female speech, and positive scores a tendency towards male speech. Tokens with values closer to 0 are gender neutral, according to the model.
@@
(“nosotras”, 0.7763) 2 (“ton”, -0.8746) 1
(“##jote”, 0.6984) 1 (“cansada” -0.8000) 1
(“Duque”, 0.6460) 1 (“Ami”, -0.6729) 1
(“Emperador”, 0.6105) 2 (“público”, -0.6722) 1
(“Majestad”, 0.6073) 3 (“desesperada”, -0.6178) 2
(“escu”, 0.6008) 1 (“contenta”, -0.5764) 1
(“tropas”, 0.5860) 1 (“Todas”, -0.4795) 16
(“infantería”, 0.5839) 1 (“esperaré”, -0.4431) 1
(“##estre”, 0.5824) 1 (“beso”, -0.4076) 1
(“encerrado”, 0.5733) 1 (“minuto”, -0.3882) 1
(“Alá”, 0.5613) 1 (“amigas”, -0.3841) 2
(“Gobernador”, 0.5448) 2 (“damas”, -0.3609) 4
(“Nevada”, 0.5178) 1 (“##ísimas”, -0.3163) 1
(“Fur”, 0.5095) 5 (“Dis”, -0.3154) 2
(“camarada”, 0.5076) 2 (“##acas”, -0.3066) 1
(“alojamiento”, 0.4962) 3 (“discursos”, -0.2966) 1
(“Fox”, 0.4954) 1 (“##món”, -0.2954) 1
(“soldado”, 0.4922) 2 (“cuerpo”, -0.2901 ) 1
(“torneo”, 0.4919) 1 (“comedia”, -0.2895) 1
(“criados”, 0.4915) 1 (“celosa”, -0.2844) 1
@@
In order to determine if there were clear themes in the tokens that the model judged as being more masculine or more feminine, we first examined the averaged attribution score of each token, calculated as the average attribution score of the token over all sentences in the validation set in which it occurs. Table 5 lists the 20 tokens deemed to be the most feminine and masculine, respectively. [^5] For convenience, the Appendix provides a translation of this table into English as Table 8.
@@
We see several patterns in the attribution scores. Included in the most feminine tokens are several word endings or words with feminine endings (cansada, desesperada, contenta, celosa, todas, ##ísimas, ##acas). These morphologically grounded indicators of female characters indicate that in Calderón’s plays, female characters either refer to themselves or other women frequently. At the semantic level, we see that some of the most feminine words have some semantic connection to interpersonal relationships between characters (amigas, beso, discursos, cuerpo, celosa). This indicates that, broadly throughout Calderón’s works, female characters are more likely to discuss themes of love and relationships than male characters. Others among these words (cansada, desesperada, contenta, celosa) also demonstrate a depiction of women as emotional – the women in these plays are on the one hand able to discuss their emotional inner worlds, but on the other hand depicted as driven or controlled by their emotions ([drouet_2017]).
This finding ties in with the results in [vsecla_2024], who show that female characters across different languages and time periods are more likely to reference their emotions than male characters. Here we can see that Calderón’s work somewhat conforms to these norms, but we will later see some evidence of the subversion of this trope in subsection 4.2.
@@
The male list contains several titles (Majestad, Emperador, Duque, Gobernador), which indicates the same pattern of more within-gender than across-gender interactions for male characters. Furthermore, these words indicate the importance of status in male-male interactions, arguably closely related to the topic of honor which was established as important across many works by Calderonian scholarship. [^6] Note, however, the absence of words for honor itself (honor, honria), arguably since honor is relevant for both men and women, even though in different contexts. Corresponding to the female indicators for conversations about relationships, the male list shows several words semantically related to combat (tropas, infantería, torneo, soldado), apparently the prototypical male topic. As previously mentioned, honor is of primary importance in Calderón’s works. In addition to the fact that war is a typically male endeavor, in the siglo de oro, male characters frequently use violence as a means to gain or restore honor ([honig_1961]). It is therefore unsurprising that these themes feature prominently as ‘masculine’ words.
@@
Curiously, nosotras is the token most attributed as male, however, it is not spoken by a male character in the corpus. It appears in the validation set 8 times, and of the 8 characters who do speak this token, 7 of them are correctly classified as female characters. Upon further investigation, it was found that nosotras appears in the training corpus, spoken by a female character that is mislabeled as male. Its listing as a female token is therefore the result of an annotation mistake, beyond the scope of our current study.
@@
We also analyzed a list of top-40 male and female tokens with regard to their morphosyntactic properties, given that word classes or word types can also be used for stylistic gender portrayal ([caliskan_2022]). We found that the most female tokens included significantly more word fragments (15), compared to the male tokens (10). Conversely, the male tokens included more nouns (24) compared to the female tokens (14). The female tokens also included slightly more adjectives (5) compared to male (3). Both used an equal number of verbs. In conclusion, there do not appear to be major differences at the level of word classes – if anything, female characters tend to use less frequent words that tend to get split up into subword tokens.
@@
Overall, this inspection of the features indicates that Calderón’s gender portrayal – within the space of features considered by the model – has two main components: (a) at the interaction level, women and men both prefer interlocutors of the same gender; (b) at the content level, women are more concerned with the domestic sphere and men the military sphere. There do not seem to be major formal linguistic differences between the two genders.
@@
Note that we only inspected the top tokens associated with either gender, and therefore this analysis cannot make any strong claim as to completeness. It is indeed likely that other topics arise from a further analysis of such tokens. Furthermore, the masking procedure we carried out during processing can be repeated after this type of analysis: If, for example, the analyst decides that tokens indicating patterns of interaction are not interesting in terms of gender portrayal, these tokens can be masked, and the model training repeated. The result is a model which should have the ‘next best’ set of features among its top-ranking gender cues.